40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 335 of 350  Not logged in ELOG logo
ID Date Author Type Category Subject
  776   Thu Jul 31 11:19:30 2008 KojiUpdateGeneralAbs. Len. Meas. ~ Resonance search trial
Last night, I tried to find the resonance of Yarm by sweeping the frequency of the injection beam.

A strong beat was present at LT_NPRO=48.7856[C_deg], the power coupling of the injection beam was estimated to be 35%. 
(Vmax_beat = 1.060[V], Vmin_beat = 0.460[V], Vno_inject = 0.664[V])

The Yarm was locked and the alignment script was executed. The PLL between the PSL beam and the injection beam was 
somehow locked.

I tried to scan the freq offset (f_PLL) at around 3.88MHz first, then at around 15.52MHz. They are supporsed to be the 
first and fourth FSR of the Yarm cavity. The Yarm transmitted power (DC) was observed to find the resonance of the 
injection beam. It would have been better to use the RF power, but so far I didnot have the RF PD prepared at the end 
transmission. I just used the DC power.

I think I saw the increase of the transmitted power by 10%, at f_PLL = 15.517 +/- 0.003 [MHz]. This corresponds to the 
arm cavity length of 38.640 +/- 0.007 [m]. The previous measurement was not so bad!

Y-arm length
e-log             length [m]
-----------------------
556(2008-Jun-24)  38.70    +/- 0.08    Cavity swinging measurement
556(2008-Jun-24)  38.67    +/- 0.03    tape & photo
This              38.640   +/- 0.007

However, I had difficulties to have more precise measurement mainly because of two reasons:
o The PLL servo is too naive, and the freqency stability of the inj beam is not enough.
  The injected beam should have the linewidth (=freq stability) narrower than the cavity linewidth.

o The PLL servo may experience change of the transfer function at around the resonance. The PLL works the other 
frequencies. However, close to the resonance, it starts to be unstable.

So the next stuffs we should do is 
o Build the PLL just using the incident beams to the ifo, not by the reflected beams.
o Build sophisticated servo to have better frequency stability.
o RF PD at the transmission.

Left the lab with Yarm locked, flipper down, shutter for the NPRO closed.
  775   Thu Jul 31 10:27:17 2008 ranaUpdatePSLPMC Scan Graphs

Quote:
Graphs of the PMC scan data that I got earlier today.

On the UNIX computers, one can use 'convert' to change these to PNG. A DC offset should be added to the transmitted
light so that the scan can be plotted with a log y-scale. And, of course, Acrobat can be used to make it into a
single PDF file.

The PMC scan always has this distortion and so the input power has to be decreased to a few mW to reduce the
thermal expansion effect; the expansion coefficient for SiO2 is ~5 x 10^-7 / K and we're worried about nm level
expansions.
  774   Thu Jul 31 10:24:32 2008 KojiUpdateGeneralIFO status
Last night I used the Y-arm for the abs length measurement. The Yarm was aligned by the script.
I left the ifo with the Yarm locked as it is the only meaningful configuration so far.
  773   Wed Jul 30 18:45:01 2008 ranaConfigurationSUSNew SUS Drift Technology
I updated the SUS DRIFT screen again, this time with a new feature.

I used Rolf's idea for the AdvLIGO status system and just made the nominals be an
unused sub-field (.SVAL) of the main INMON record. Then I wrote a .csh script to
use tdsavg to average the current INMON vals and insert that as the .SVAL. The next
script should read the .SVAL and set the HIHI and LOLO based on this.

Of course, the values I have just entered are no good because our suspensions are in
a bad state but we can run this script (SUS/setDriftNoms) next time things are good.

And...even better would be if someone were to do the same but used mDV to grab the
minute trend in the past good times instead of the tdsavg (which can't go in the past).
  772   Wed Jul 30 16:35:56 2008 EricUpdatePSLPMC Scan Graphs
Graphs of the PMC scan data that I got earlier today.

PMCLongScanWide.tiff shows the transmission intensity and PZT voltage plotted against time for a longer scan of the PMC (~120 seconds for one sweep).

PMCLongScanPeak.tiff is the same scan zoomed in on the primary peak. This scan was done with the laser power at around 1/3 its original value. However, scans done at around 1/6 the original value have peaks that are just as messy.

PMCShortScanWide.tiff shows the intensity and voltage for a more rapid scan (~30 second for one sweep). The black lines show how the peak positions are at very different PZT voltages (a difference of ~10 volts in both cases).

PMCShortScanPeak.tiff is zoomed in on the primary peak. The peak is much cleaner than for the long scan (less time for the laser's heat to expand the mirror?), though it is likely still too messy to reliably fit to a lorentzian.
Attachment 1: PMCLongScanPeak.tiff
Attachment 2: PMCLongScanWide.tiff
Attachment 3: PMCShortScanPeak.tiff
Attachment 4: PMCShortScanWide.tiff
  771   Wed Jul 30 15:28:08 2008 robUpdateLSCY arm locked

By using a combination of the SUS-DRIFT mon screen and the optical levers (which turned out pretty well) I steered the BS, ITMY, and ETMY back to their previous positions, and was able to lock the Y arm. The "Restore Y Arm" script on the IFO_CONFIGURE screen works. I couldn't test the alignment script, as a dump truck/construction vehicle showed up and started unlocking the MC.
  770   Wed Jul 30 15:12:08 2008 ranaSummaryIOOHistory of the MC abs length
> I was notified by Rob and Rana that there were many measurements of the MC abs length (i.e. modulation
> frequencies for the IFO.) between 2002 and now.

I will just add that I think that the Marconi/IFR has always been off by ~150-200 Hz
in that the frequency measured by the GPS locked frequency counter is different from
what's reported by the Marconi's front panel. We should, in the future, clearly indicate
which display is being used.
  769   Wed Jul 30 13:52:41 2008 EricSummaryCamerasWeekly Summary
I tracked the tendency for ezcaPut to fail and sometimes seg-fault in the camera code to a conflict between the camera API and ezca, either on the 
network level or the thread level.  Since neither are sophisticated enough to provide controls over how they handle these two things, I instead 
separated the call to ezcaPut out into a small, separate script (a stripped down ezcawrite), which the camera code calls at the system level.  This is a 
bit hacky of a solution, but its the only thing that seems to work.

I've developed a transformation based on Euler angles that should be able to take the 4 OSEMs in a picture of the end mirror and use their relative 
positions to determine the angle of the camera to the optic.  This would allow the position data determined by the fitting software to be converted 
from pixels to meaningful lengths, and should aid any servo-ing done on the beams position.  I've yet to actually test if the equations work, though.

The servo code needs to have slew rate limiters and maximums/minimums to protect the mirrors written in to it before it can be tested again, but I 
have no idea what reasonable values for these limits are.

Joe and I recently scanned the PMC by driving C1:PSL-PMC_RAMP with the trianglewave script over a range of -3.5 to -1.25 (around 50 to 150 volts 
to the PZT) and read out C1:PSL-ISS_INMONPD to measure the transmission intensity.  This included slightly under 2 FSRs.  For slow scans (covering 
the range in 150 to 300 s), the peaks were very messy (even with the laser power at 1/6 its normal value), and it was difficult to place where the 
actual peak center occurred.  For faster sans (covering the range in 30 seconds or so), the peaks were very clean and nearly symmetric, but were 
not placed logically (the same peak showed up at two very different values for the PZT voltage in two separate runs).  I don't have time to put 
together graphs of the scans at the moment; I'll have that up sometime this afternoon.
  768   Wed Jul 30 13:14:03 2008 KojiSummaryIOOHistory of the MC abs length
I was notified by Rob and Rana that there were many measurements of the MC abs length (i.e. modulation 
frequencies for the IFO.) between 2002 and now.

So, I dig the new and old e-logs and collected the measured values of the MC length, as shown below. 

I checked the presence of the vent for two big steps in the MC length. Each actually has a vent. 
The elog said that the tilt of the table was changed at the OMC installation in 2006 Oct.
It is told that the MC mirrors were moved a lot during the vent in 2007 Nov.

Note:
o The current modulation freq setting is the highest ever.
o Rob commented that the Marconi may drift in a long time.
o Apparently we need another measurement as we had the big earthquake.

My curiosity is now satified so far.

Local Time	3xFSR[MHz]	5xFSR[MHz]	MC round trip[m]	Measured by
----------------------------------------------------------------------------
2002/09/12	33.195400 	165.977000 	27.09343		Osamu
2002/10/16	33.194871 	165.974355 	27.09387		Osamu
2003/10/10	33.194929 	165.974645 	27.09382		Osamu
2004/12/14	33.194609 	165.973045 	27.09408		Osamu
2005/02/11	33.195123 	165.975615 	27.09366		Osamu
2005/02/14	33.195152 	165.975760 	27.09364		Osamu
2006/08/08	33.194700 	165.973500 	27.09401		Sam
2006/09/07	33.194490 	165.972450 	27.09418		Sam/Rana
2006/09/08	33.194550 	165.972750 	27.09413		Sam/Rana
----2006/10 VENT OMC installation	
2006/10/26	33.192985 	165.964925 	27.09541		Kirk/Sam
2006/10/27	33.192955 	165.964775 	27.09543		Kirk/Sam
2007/01/17	33.192833 	165.964165 	27.09553		Tobin/Kirk
2007/08/29	33.192120 	165.960600 	27.09611		Keita/Andrey/Rana
----2007/11 VENT Cleaning of the MC mirrors
2007/11/06	33.195439 	165.977195 	27.09340		Rob/Tobin
2008/07/29	33.196629 	165.983145 	27.09243		Rob/Yoichi
Attachment 1: MC_length.png
MC_length.png
  767   Wed Jul 30 13:09:40 2008 josephb, EricConfigurationPSLPMC scan experiment
We turned the PSL power down by a factor of 4, blocked one half of the Mach Zehnder and scanned the PMC by applying a ramp signal to PMC PZT. Eric will adding plots later today of those results.

We returned the power to close to original level and removed the block on the Mach Zehnder, and then relocked the PMC.
  766   Wed Jul 30 13:08:44 2008 Max JonesUpdateComputer Scripts / ProgramsWeekly Summary
This week I've been working on the noise budget script. The goal is to add Siesmic, Darm, Mich, Prc and magnetometer noise. I believe I've added Seismic noise in a reasonable and 40m specific manner (please see the attached graph). The seismic noise in the noise budget at 100 Hz was 10 times higher than that predicted by Rana in elog #718. This could be due to the fact that graph is taken from data today when the device is unlocked and construction workers are busy next door. I am currently trying to fix the getDarm.m file to add the DARM source to the noise budget. I have run into several problems, the most pressing of which was that the C1:LSC-DARM_ERR channel is zero except with the interferometer is being locked. According to Rob, we only save data for approximately a day (we save trends for much longer but this is insufficient for the noise budget script) and sometimes we are not locked the night before. Rob showed me how I may introduce an artificial noise in the DARM_ERR signal but I'm having trouble making the script output a graphic. I'm still unsure how to make the getDarm function 40m specific.

Today I will start working on my second progress report and abstract.
Attachment 1: C1_NoiseBudgetPlot.pdf
C1_NoiseBudgetPlot.pdf
  765   Wed Jul 30 12:36:19 2008 SharonUpdate Weekly update
This week included many computer's issues. I tested Alex's new C code (the one that saves the FIR coefficients and restores them when you start running the code again). Seems there is an improvement in the adaptation time, but not a significant one (more details on the coming report). I had to recompile simulink and the FB whenever I wanted to find a solution for taking the record of those coefficients. This is so I could simulate the adaptive filter with a regular IIR filter and compare the two.
After Rob tried to help and it seems to be an impossible to a huge hassle mission, we thought of a different method to do this. I re-compiled the old simulink file and restored the .ini file and all should be back in place. Instead of finding the FIR coefficients, I am going to use one noise source in the adaptive filter, stop the adaptation (by setting mu and tau to 0), and put excitation instead of the noise signal. The transfer function I will get between the exc. and MC1_IN1 is the filter I am looking for.

Also seems that whenever I get the MC unstable, and the adaptive code stops itself, it doesn't come back. Setting the reset flag to a different number (anything other than 0) and pressing the reset button will get it working again, but the CPU will always flip and the ASS computer needed a restart. Still haven't found a problem in the C code, but that's the plan. Moreover, I want to change Alex's code, so that instead of starting from zero like in Matt's code, or starting from the old coefficients like in Alex's, it is going to calculate a Wiener Filter as the first set of coefficients. This will hopefully reduce the adaptation time.

I have also been working on my progress report, and stood in line for the MC... Still standing...
  764   Wed Jul 30 12:03:44 2008 MashaSummaryAuxiliary lockingweekly summary
I've been learning about mode matching/beam propagation, so I can work on getting more
light into the fiber and increase the phase noise signal. I am also looking into phase
lock loops and noise in the fiber stabilization system to understand the noise sources
and figure out what our goals are in fiber stabilization.

In the lab, I've reproduced the Mach Zehnder interferometer that I had at the 40m, now
with a 50m fiber in one arm. I have done some preliminary fiber noise measurements
and revised estimations of noise sources (see attached plots). Once the digital
acquisition system is back up, I will be able to better manipulate the signals to cancel
laser amplitude noise and amplitude noise from variation in the amount of coupling into
the fiber. Some improvements in progress are more stable mounts for the fiber couplings,
faraday isolator, and better mode matching with the fiber.

Also working on my progress report.
Attachment 1: mz_fiber_noise0730.png
mz_fiber_noise0730.png
Attachment 2: setup0730.JPG
setup0730.JPG
Attachment 3: MZnoise_sources0727.png
MZnoise_sources0727.png
  763   Wed Jul 30 01:08:50 2008 ranaSummarySUSSUS Drift Screen
This is a snap of the SUS Drift screen with all of the optics biases set back to their nominal
values except for the MC which Rob aligned and I didn't feel like mis-aligning. The reference
on the screen is from 3/25 when Andrey felt that Rob had a good IFO alignment.

Anything more than a few thousand is significant and more than 10k means something is wrong:



I wailed on the PRM for awhile and was able to loosen it up a little. The LL & LR sensors now
show some life one the dataviewer. The UL & UR are still railed ~1.6 V so that means that the
optic is pitched back. With aggressive pitch wailing I can see the PRM's ULR/UR sensors go
rail to rail so that means that the magnets are still on - although they may be half busted.
If they're OK we should be able to just re-sling this guy.


Did the same on SRM. The OSEM values have shifted on these, but not disastorously. The SIDE,
however, is completely unresponsive. The little signal I see when driving is is probably just
capacitive pickup in the cables. Have to vent to fix this one.


ITMX Has good life in all but the LR & UR channels. They respond, but the signal is very weak.
Seems like these magnets have not fallen off but that they are not between the LED/PD anymore.


ITMY seems ok. Check the spectra to be sure.


BS seems ok as well. Swings freely and no kinks in the swinging sensor waveforms. Check the spectra.
Attachment 1: infection-2.png
infection-2.png
  762   Wed Jul 30 00:42:04 2008 ranaUpdateSUSTrends and file formats
I propose that we do not use .eps format but .pdf instead. For images like the plots Sharon
has below we should use only .png and for pictures like what Steve posted, use JPG or PNG.

PDF is a standard and light weight. PNG is very good for plots/lines and is lossless. JPG does
a good job with regular camera pictures because we don't really care about the compression
loss on those.

Here's a trend of the UL sensors for all the optics - conversion is 32768 cts / mm. You can see
that the quake was just before 19:00 UTC (noon our time). The events an hour after are when
Rob, Jenne, and I start exciting the optics to shake them loose - wanging the pit/yaw sliders
around is not violent enough and so I injected a 130000 count sine wave at 0.5 Hz so as to
create a high force square wave. This seems to have worked for ETMY but no such luck yet with
the others.
Attachment 1: Untitled.png
Untitled.png
  761   Tue Jul 29 23:04:34 2008 YoichiUpdatePSLFSS loop transfer functions

Quote:

The measurement of the PZT open loop TF is very suspicious. According to this, the PC path has a very large gain even at very low frequencies (there is no cross over above 1kHz). This cannot be true. Maybe the cavity's optical gain was low when it was locked with only the PZT. I will re-measure it.


I measured it again and found that the loop was oscillating at 13.5kHz. I think this oscillation prevented the ref. cavity from building up the power and consequently lowered the optical gain making it marginally stable. So the PZT path open loop TF posted in the previous entry is wrong.

I was able to stop the oscillation by lowering the gain down to CG=-7.6dB and FG=-8.78dB.
The first attachment shows the measured open loop transfer function.
Since the gain setting is different from when the over all open loop TF was measured, I scaled the gain (attachment 2).
However, this plot seems to have too much gain. Scaling it down by 20dB makes it overlap with the over all open loop TF.
Maybe the gain reading on the EPICS screen is wrong. I will measure the actual gain tomorrow.
Attachment 1: OpltfPZTOnlyRaw.eps
OpltfPZTOnlyRaw.eps
Attachment 2: OpltfPZTOnly.eps
OpltfPZTOnly.eps
  760   Tue Jul 29 21:04:55 2008 SharonUpdate OSEM's Power Spectrum
From 16:30 this afternoon
Attachment 1: ITMY2.JPEG
ITMY2.JPEG
Attachment 2: ITMY.JPEG
ITMY.JPEG
Attachment 3: ITMX2.JPEG
ITMX2.JPEG
Attachment 4: ITMX.JPEG
ITMX.JPEG
Attachment 5: ETMY2.JPEG
ETMY2.JPEG
Attachment 6: ETMY.JPEG
ETMY.JPEG
Attachment 7: ETMX2.JPEG
ETMX2.JPEG
Attachment 8: ETMX.JPEG
ETMX.JPEG
  759   Tue Jul 29 19:53:19 2008 KojiUpdateSUSPRM photos from the south window
Steve and Koji

We took some photos of PRM from the south window.
You can see one of the side magnets, a wire stand-off, and the wire itself from the round hole.
So, the wire looks OK.

For the coils, we could see only one coil. The magnet is apparently too high.
Attachment 1: PRM_from_South_Window1.jpg
PRM_from_South_Window1.jpg
Attachment 2: PRM_from_South_Window2.jpg
PRM_from_South_Window2.jpg
  758   Tue Jul 29 19:41:38 2008 YoichiUpdatePSLFSS loop transfer functions
Last night I measured a bunch of transfer functions on the FSS loop.
All the loop gains were measured with the common gain = 30db and the fast gain = 18dB.

(1) The first attachment is the overall open loop transfer function of the FSS loop. I put a signal from the Test IN2 and observed signals from IN1 and IN2.
The UGF is about 180kHz.
By increasing the RF amplitude going to the EOM (i.e. increasing the sideband power), I can further increase the gain of the servo.
However, it made the PC drive immediately crazy. Probably it was some oscillation.

(2) Then I locked the ref. cav. with only the PZT actuator. I did so by simply unplugging the cable going to the PC.
Surprisingly, the cavity locked with the *same* gain setting as before. The second attachment shows the open loop transfer function measured in this configuration. It seems wrong, I mean, it should be unstable. But the cavity locked. A mystery.

(3) The third plot is the measured TF from the Test IN1 of the FSS board to the fast out (output to the PZT).

(4) By dividing the TF measured in (2) with the TF of (3), I got the response of the PZT times the cavity response. This is shown in the attachment 4.

(5) We can guess the open loop TF of the PC path by subtracting the TF in (2) from (1). It is shown in the attachment 5.

(6) The filter shape of the PC path is measured by injecting signal from the Test IN1 of the FSS board and observing it at the PC output. Since it is a high voltage output, I reduced the common gain to -8.5dB during the measurement. The attachment 6 is the measured filter shape. The gain is corrected to show what it should look like when the common gain = 30dB.

(7) By dividing (5) with (6), I plotted the response of the PC times the cavity response in the attachment 7. Since the 1/f cavity pole and the response of the PC, which is proportional to f, should cancel out, we expect a flat response above the cavity pole frequency (38kHz). You could say it is a sort of flat, if you have obscured eyes.

The measurement of the PZT open loop TF is very suspicious. According to this, the PC path has a very large gain even at very low frequencies (there is no cross over above 1kHz). This cannot be true. Maybe the cavity's optical gain was low when it was locked with only the PZT. I will re-measure it.
The plot (4) is also strange becaues it does not show the low pass feature expected from the cavity pole.
Attachment 1: OverallOPLTF.eps
OverallOPLTF.eps
Attachment 2: OpltfPZTOnly.eps
OpltfPZTOnly.eps
Attachment 3: PZTFilter.eps
PZTFilter.eps
Attachment 4: PZTxCavityPole.eps
PZTxCavityPole.eps
Attachment 5: OpltfPCOnly.eps
OpltfPCOnly.eps
Attachment 6: PCFilter.eps
PCFilter.eps
Attachment 7: PCxCavityPole.eps
PCxCavityPole.eps
  757   Tue Jul 29 18:15:36 2008 robUpdateIOOMC locked

I used the SUS DRIFT MON screen to return the MC suspensions to near their pre-quake values. This required fairly large steps in the angle biases. Once I returned to the printed values on the DRIFT screen (from 3/08), I could see HOM flashes in the MC. It was then pretty easy to get back to a good alignment and get the MC locked.
  756   Tue Jul 29 14:38:02 2008 robUpdateSUSETMY and PRM have EQ related problems

Quote:
The attached trend shows that ETMY and PRM both had large steps in their sensors
around the time of the EQ and didn't return afterwards. The calibration of the
OSEM sensors is ~0.5 mm/V. The PRM sensors respond when we give it huge biases
but there is very little change in the ETMY. Almost certainly true that the
optics have shifted in their wire slings and that we will have to vent to
examine and repair at least ETMY.

Jenne is looking at the spectra of the other suspensions to see if there is
other more subtle issues.


Some additional notes/update:

ETMY, PRM, & MC2 had OSEM signals at a rail (indicating stuck optics). Driving the optics with full scale DAC output freed ETMY and MC2, so while these may have shifted in their slings it may be possible to avoid a repair vent. PRM is still stuck. One OSEM appears to respond with full range to large drives, but the other three face OSEMS remain disturbingly near the rail (HIGH, which is what would happen if a magnet fell off).
  755   Tue Jul 29 13:54:08 2008 ranaUpdateSUSETMY and PRM have EQ related problems
The attached trend shows that ETMY and PRM both had large steps in their sensors
around the time of the EQ and didn't return afterwards. The calibration of the
OSEM sensors is ~0.5 mm/V. The PRM sensors respond when we give it huge biases
but there is very little change in the ETMY. Almost certainly true that the
optics have shifted in their wire slings and that we will have to vent to
examine and repair at least ETMY.

Jenne is looking at the spectra of the other suspensions to see if there is
other more subtle issues.
Attachment 1: Untitled.png
Untitled.png
  754   Tue Jul 29 11:50:01 2008 JenneUpdateEnvironment5.6 Earthquake
Earthquake Details
Magnitude 5.6
Date-Time

* Tuesday, July 29, 2008 at 18:42:15 UTC
* Tuesday, July 29, 2008 at 11:42:15 AM at epicenter

Location 33.959°N, 117.752°W
Depth 12.3 km (7.6 miles)
Region GREATER LOS ANGELES AREA, CALIFORNIA
Distances

* 3 km (2 miles) SW (235°) from Chino Hills, CA
* 8 km (5 miles) SE (127°) from Diamond Bar, CA
* 9 km (5 miles) NNE (23°) from Yorba Linda, CA
* 11 km (7 miles) S (178°) from Pomona, CA
* 47 km (29 miles) ESE (103°) from Los Angeles Civic Center, CA

Location Uncertainty horizontal +/- 0.3 km (0.2 miles); depth +/- 1.3 km (0.8 miles)
Parameters Nph=144, Dmin=8 km, Rmss=0.42 sec, Gp= 18°,
M-type=local magnitude (ML), Version=1
Source

* California Integrated Seismic Net:
* USGS Caltech CGS UCB UCSD UNR

Event ID ci14383980

All the watchdogs tripped. I'll put them back after lunch, after the optics have had time to settle down.
  753   Tue Jul 29 09:12:43 2008 KojiConfigurationIOOMC length measurement
I found that the prev modulation freq had been determined with a same kind of measurement by Osamu, which also looked accurate.
http://www.ldas-sw.ligo.caltech.edu/ilog/pub/ilog.cgi?group=40m&task=view&date_to_view=09/12/2002&anchor_to_scroll_to=2002:09:12:17:10:30-ajw

(There is also a document by Dennis to note about this measurement
http://www.ligo.caltech.edu/docs/T/T020147-00.pdf )

So, it means that the round trip length of the MC shortened by 1mm in the 6 years.
New              OLD
--------------------------
27.0924          27.0934    [m]

Quote:
rob, yoichi

We measured the length of the mode cleaner tonight, using a variant of the Sigg-Frolov method.
....
The new f2 frequency is:
New              OLD
--------------------------
165983145        165977195
  752   Tue Jul 29 01:03:17 2008 robConfigurationIOOMC length measurement
rob, yoichi

We measured the length of the mode cleaner tonight, using a variant of the Sigg-Frolov method. We used c1omc DAC outputs to inject a signal (at 2023Hz) into the AO path of the mode cleaner and another at DC into the EXT MOD input of the 166MHz IFR2023A. We then moved an offset slider to change the 166MHz modulation frequency until we could not see the 2023Hz excitation in a single-bounce REFL166. This technique could actually be taken a step further if we were really cool--we could actually demodulate the signal at 2023Hz and look for a zero crossing rather than just a powerspec minimum. In any case, we set the frequency on the Marconi by looking at the frequency counter when the Marconi setting+EXT MOD input were correct, then changed the Marconi frequency to be within a couple of Hz of that reading after removing the EXT MOD input. We then did some arithmetic to set the other Marconis.

The new f2 frequency is:

New              OLD
--------------------------
165983145        165977195

  751   Mon Jul 28 23:41:07 2008 robConfigurationPSLFSS/MC gains twiddled

I found the FSS and MC gain settings in a weird state. The FSS was showing excess PC drive and the MC wouldn't lock--even when it did, the boost stage would pull it off resonance. I adjusted the nominal FSS gains and edited the mcup and mcdown scripts. The FSS common gain goes to 30dB, Fast gain to 22dB, and MCL gain goes to 1 (which puts the crossover back around ~85 degrees where phase rises above 40 degrees).
  750   Mon Jul 28 17:58:05 2008 SharonUpdate TOP screen changes
I wanted to test the adaptive code with a downsampling rate of 32 instead of 16. To do this I entered a 32 Hz ((2048/32)/2 - match Nyquist Freq.) low pass filter on the ERROR EMPH, MC1 and the relevant PEM channels.
  749   Mon Jul 28 17:44:07 2008 ranaUpdatePSLPMC PZT v. temperature
This plot shows that the PMC PZT has ~20 Vpp fluctuations on a 24 hour timescale
which is correlated to the 24 hour temperature fluctuations. By contrast, the MZ
has ~75 Vpp
.
Attachment 1: Untitled.pdf
Untitled.pdf
  748   Mon Jul 28 15:54:04 2008 KojiUpdateGeneralAbs. Len. Meas. ~ More on the beat / the PLL setup
Alberto and Koji,

Last Friday evening, Koji found that the power adj setting (indicated by ADJ) of the NPRO was somehow set to be 
ADJ=-45 and yielded the output power of about 200mW instead of 700mW. This is not good because too small pump power 
varies thermal conditions of the crystal such as thermal lensing, thermal gradient, and os on. The ADJ setting and the 
crystal temperature had been restored to ADJ=0 and LT=~48deg (nominal of the controller), respectively.

Today we tried the quest of the beating again and the above power setting helped a lot! The beating was immediately 
found at LT=48.55deg that is very close to the laser's nominal temp. Also the beating got significantly bigger. 
After the alignment adjustment 50%-intenisity modulated signal was obtained. From the power calculation it was 
estimated that the power coupling of the injected beam is to be 12%~13%. This not so good yet, but something which we 
can work.

This time the modulation structure of the PSL beam was clearly observed. I could obtain the beating of the injection 
beam with the carrier, the upper/lower sidebands of the 33MHz and 166MHz modulations, and the 2nd order of the 
33MHz. They were beautiful as if working with an OSA. Very nice.

In reality, those additional intenisty modulations as well as the residual 33MHz signal from the main IFO are 
disturbing for the PLL to be locked at the proper frequency. So, now Alberto is working on a passive LPF with 
notch at 33MHz. The design was already done. This allows us to work up to 20MHz and at the same time, provides 
60dB attenuation at 33MHz (in principle). Very cool.

Koji, on the other hand, continued to work with the PLL servo with some ready-made passive filters. Owing to the 
fillters, the error signal was cleaner and the PLL was locked at the proper frequency. The PLL setup is as attatched. 
Sideband rejection filter will be replaced to Alberto's one. The photo is the display of the RF spectrum analyzer with 
beat locked at 8MHz.

So the next step, we try to find the resonances of the arm cavity with the injection beam once the IFO comes back.

At the last of the experiment "Last autoalignment" was restored, the flipper for the 
inj beam was down, and the shutter for the NPRO was closed.
Attachment 1: PLL_setup.png
PLL_setup.png
Attachment 2: beat_at_8MHz.jpg
beat_at_8MHz.jpg
  747   Mon Jul 28 12:02:32 2008 SharonUpdate accelerometers settings
Jenne, Sharon


We looked again at the channels of the accelerometers and there are some updates. Last time when we reported, we crossed the ADAP channels and the accelerometer. Now that there is a new MEDM screen, with which you can control which channels goes to which adaptive channels, this has no meaning...
Therefore, the channels that go with the noise source channels are:

PEM 15 MC1_X
PEM 16 MC1_Y
PEM 17 MC1_Z
PEM 18 MC2_X
PEM 19 MC2_Y
PEM 20 MC2_Z
PEM 21 SEIS

disregard the last post regarding these channels by Jenne, since I am changing the ADAP channels all the time...
  746   Mon Jul 28 11:20:13 2008 JenneUpdatePSLWork on the FSS and Reference Cavity
[Yoichi, Jenne, Koji]

The Reference Cavity's saga continues....

Thursday, Yoichi and I worked to change the beam that we chose from the 2nd pass through the AOM, to the first order beam rather than the 2nd order beam (see elog #726). After choosing the correct beam, we get 29mW incident on the reference cavity (compared with 4mW before any work began). We adjusted the angle of the AOM in the plane of the table, and got up to 30.6mW. We adjusted the tip/tilt of the AOM and got to 30.7mW (the tip/tilt adjustment made a more significant difference in the work described in elog #726, but after that work, it was probably already pretty close to optimized). We noticed that for the above measurements, we had 2 beams through the Polarizing Beam Splitter and Waveplate (one very dim), so after excluding that beam, the power meter read 30.4mW. We adjusted the curved mirror a little, and got 30.8mW incident on the reference cavity.

We then put a triangle wave into the offset of the MC Servo Board using the "trianglewave <channel> <center> <amplitude> <period> <runtime>" command in a terminal screen. This changes the voltage to the VCO, and thus the frequency response of the AOM. We watch the diffracted spots from the second pass through the AOM, and confirm that the beam we have chosen is not moving, and all the others are. By symmetry, if we chose the first order beam after the first pass through the AOM, and then again chose the first order beam after the second pass, the resulting beam will not move with the frequency change of the AOM.

We saw 1.50V (Refl. PD, unlocked) on the 'scope after aligning the optics to make the newly chosen beam hit the input mirror of the reference cavity. Order of operations for this alignment:
  • Recenter the beam on the 2 lenses that are just after the PBS and the waveplate
  • Adjust pitch and yaw of the two steering mirrors until the beam reflected off the input mirror of the reference cavity is parallel to the incident beam
    • Use a sensor card to check the alignment of the incident and reflected beams, and adjust the steering mirrors to get the alignment close
      • Note the amplitude of the DC output of the Refl. PD with the iris completely open. Close the iris until the signal decreases by ~50%, then adjust the steering mirrors until the original amplitude is regained. Repeat until the iris can be almost completely closed but the Refl. PD signal doesn't change
    • Watch the DC output of the Refl. PD, and maximize the signal on a 'scope
    • Sweep the PZT of the laser using a function generator into the RAMP input on the FSS board (~10Vpp at ~1Hz), OR sweep the temperature of the laser using the trianglewave function on the SLOW FSS channel (amplitude~0.5, period~50)
    • Watch the modes that resonate in the cavity, and adjust pitch and yaw of the steering mirrors to get closer to the TEM00 mode
    • When the TEM00 mode appears in the sweep, stop the sweep, and lock the cavity
    • Watch the DC output of the Transmitted PD, and maximize the signal on a 'scope
  • Celebrate!

After all of this adjusting,
Refl. PD (unlocked) = 1.48V
Refl. PD (locked) = 680mV
Trans. PD (locked) = 6.28V
Power reflected (unlocked) = 26.28mW
Power transmitted (locked) = 13.89mW
Thus, 53% transmission

Next: check the amount of power transmitted by reducing the amplitude of the RF modulator. This reduces the amount of power used by the sidebands, and so should increase the transmission.
Power incident = 27mW
Power transmitted = 17.2mW
Thus, 64% transmission
We then put the RF modulator back where it was originally.

We then replaced the lens mounts for the f=802 and f=687 lenses between the AOM and the reference cavity, to the new mounts that Yoichi bought. Koji helped me realign into the reference cavity, and we got:
Refl. PD (unlocked) = 1.48V
Refl. PD (locked) = 880mV
Trans PD (locked) = 4.64V
Power incident = 26.97mW
Power transmitted = 10.39mW
39% transmission
Since more mode matching etc. is in the works, we left this for the night.

On Friday, we changed the setup of the cameras and PDs for both reflection and transmission, to avoid saturating the PDs and cameras.

On the Refl. side of the reference cavity, we put a W2-PW-1025-UV-1064-45P pickoff between the last mirror and lens before the camera and PD. We moved the camera to the pickoff side of the new optic. We then replaced teh 45UNP beam splitter that split the beam between the PD and the camera with a Y1-1037-45P highly reflective mirror, and put the PD in the old camera location.

On the Trans. side of the ref. cavity, we replaced the BSI-1064-50-1025-45S with a W2 pickoff, and replaced the Y1-1037-45-P highly reflective mirror with the 50/50 beam splitter that was replaced by the W2.

Now we have:
Refl. PD (unlocked) = 1.68V
Refl. PD (locked) = 640mV
Trans PD (locked) = 4.24V
Power incident = 25mW
Power transmitted = 14.48mW
58% transmission

Koji pointed out that when remounting, I had put the f=802 lens ~2cm away from its original position (along the z-axis), so I moved the lens back to where it should be, and realigned into the reference cavity. Since Rana was working on the PMC at the same time, the laser was turned down by about a factor of 100, so my starting measurements were:
Refl. PD (unlocked) = 23.6mV
Refl. PD (locked) = 10.2mV
Trans PD (locked) = 56mV
Power incident = 0.35mW
Power transmitted = 0.16mW
46% transmission

Since it was late on Friday by the time everything was realigned into the ref. cavity (I'm still working on my optics aligning skills), I forgot to measure the transmission after all of my work. I'll do that today (Monday) as soon as Sharon/Koji are done working with the IFO this morning. Also, I'll put up before/after pictures as soon as I find the camera...it seems to have walked off.

UPDATE:
Ref. Cav. measurements after Friday's alignment (and after turning the laser power back up to normal):
Refl. PD (unlocked) = 1.58V
Refl. PD (locked) = 304mV
Trans PD (locked) = 3.68V
Power incident = 24.96mW
Power transmitted = 16.45mW
66% transmission


To do: Start the actual mode-matching into the reference cavity.
  745   Sun Jul 27 23:06:17 2008 ranaUpdatePSLPMC, MZ, MC-MMT, etc.
With the new PMC now in I aligned the MZ to the new beam (there is sadly no steering
between the PMC and the MZ).

I also removed the pickoff that we had put before the MZ just in case we wanted to
move the FSS pickoff to there - its been 2 years now so I guess its not going to happen.


The new PMC's cavity axis seems to be a few hundred microns higher than the old one. So I
tried to move the MZ EOMs to compensate but ended up also steering all of the MZ's mirrors
to get the contrast good, the beam onto the ISS PDs, centered (sort of) onto the MMT lenses
and onto the periscope.

Along the way I also removed some of the vestigial squeezer stuff around the power control
PBS. The output of the PBS now goes directly into the high power dump with no steering. This
eliminated around a dozen clamps, bases, etc. and a couple of mirrors.


The MC is locked on the low power beam we have running through everything. I restored the
PSL launch beam just using the MC-WFS and it locked on a TEM00. So now we know that we
really don't need the PSL quads for this as long as the MC1 angle is stable.

The good news is that the PMC PZT voltage is now flat: the problem must have really been with
the PZT
and not the cabling or notch box like I had wondered about.

Todo:
-----
1) Continue mode matching into the PMC. Its transmission now is around the same as the
   old one.

2) Put a UHV foil covered lead brick onto the PMC to quiet it down.

3) Characterize the PMC loop and retune the body notch for the new body.

4) Tweak the MZ alignment to minimize the RFAM. We can use StochMon to do this as
   long as we have the MC WFS turned off or we can put in a flipper to take the
   beam before the MC and send it to the StochMon RFPD.

5) Re-align onto the ISS.

6) Install irises around the periscope for the beam. The old iris there is way off.

7) Fix PSL ANG and center both POS and ANG.
  744   Sun Jul 27 20:49:21 2008 ranaConfigurationComputersNTP
After Aidan did whatever he did on op440m, I had to restart ntpd. I noticed it didn't actually do
anything so I restarted it by hand with the '-l' option to make a logfile. Essentially, the
problem is that NTPD is not allowed access to the outside world's NTP servers by our NAT router;
this should be fixed.

So for now I set all of the .conf files to point to rana and nodus' IP addresses. According to the
log files, that is successful. Rosalba and Mafalda, however, seem to have correct time but are
looking at rhel.ntp.pool.org and time.nist.gov, respectively. Maybe these have special rules?

For reference, the linux machines' conf files are /etc/ntp.conf
and the solaris machines' conf files are /etc/inet/ntp.conf

I also logged into dcuepics (aka scipe25) and did as instructed.
  743   Sun Jul 27 20:25:49 2008 ranaConfigurationEnvironmentOffice Temperature increased to 75 F
Since we have the chiller for the PSL chiller now, I've just increased the office area
temperature set point by 2 F
to 75 F to see if the laser will still behave.
  742   Sat Jul 26 15:09:57 2008 AidanUpdateComputersReboot of op440m

I was reviewing the PSL Overview screen this afternoon and op440m completely froze when I center-clicked on the REF CAVITY TRANSMISSION indicator. It was unresponsive to any keyboard or mouse control. The moon button had no effect to shut the machine down.

Called Alberto in and we logged into op440m from rosalba. From there we logged in as 'root' and run a shutdown script '/usr/sbin/shutdown -i S -g 1'. The medm screens started disappearing from the op440m display and we were eventually asked to enter System Maintenance Mode. From here we selected RUN LEVEL 5: "state 5: Shut the machine down so that it is safe to remove the power". Following this the machine turned itself off.

We powered it back on, logged back in as controls and restarted the medm screens. Everything seems to be running fine now.
Aidan.
  741   Fri Jul 25 19:57:18 2008 JenneUpdatePSLRef Cav & PMC
"PMC is in, but is still being worked on. Leave it alone." ---Rana

Ref. Cavity is locked again. Still a work in progress. I think we're ready to mode match on Monday. ---Jenne
  740   Fri Jul 25 17:32:46 2008 SharonUpdate ASS computer
So, it seems a bit too complicated getting the coefficients the way I wanted it to happen (simulink-.ini...).
I returned everything to the way it was and it's all working. The new plan is to choose the specific channel I want to find its instantanous coefficients, let the adaptive code run for a while, setting mu and tau to zero (freezing the coefficients), and exciting the noise signal channel taking the transfer function. This way I can find the filter I want to simulate with an IIR filter.
Once I have the mode cleaner to myself, I'll start posting results.
  739   Fri Jul 25 13:30:53 2008 SharonUpdate Changes in ASS computer
I editted the simulink diagram of the ASS computer so it now has 2 more channels reading 2 sets of the FIR coefficients to match Alex's changes in the C code.
The new simulink has already been compiled and can be found in /cvs/cds/caltech/users/alex/cds/advLigo/src/epics/simLink/ass.mdl
I backed up the old file and it's also in that folder under ass_BAK_24_jul.mdl

There is also a backup of the old ASS.ini file in caltech/chans/daq/C1ASS_BAK_24_jul.ini

Will update once it's all set and running
  738   Fri Jul 25 10:48:13 2008 KojiUpdateGeneralAbs. Len. Meas. ~ Realignment / beating / PLL trial
Alberto and Koji

o We worked for the abs length measurement setup on Thursday night.
o At the last of the work Koji left the 40m lab at 2AM. "Last autoalignment" was restored. The flipper for the 
inj beam was down. The shutter for the NPRO was closed.

----
o The alignment of the injection beam (NPRO) was re-adjusted.
o The laser crystal temp (LT) of the NPRO was scanned.
o After a long struggle the beat was found at about LT=61deg(!). I think this is almost at the maximum temp 
for the NPRO. Note that this is not the diode temp, and therefore it will not damage the laser as far as the 
TEC for the crystal works.

o Only the X arm was aligned.
o The alignment of the injection beam was adjusted such that the beating amplitude got maximum.
o At the faraday of the NPRO, we had 2.4V_DC and 1.8V_DC with and without the inj beam, respectively. The 
beating amplitude was 200mVpp (at around 2.4V).
o With a simple calculation, the mode overlapping of tghe injection beam is only 0.0023. Ahhh. It is too weak. 
In the modematching or something must be wrong. 
o The position of the mode matching lens was tweaked a little. It did not help to increase the beat ampitude. 
Even worse. (The lens was restored and the values above was obatined with the latest setting.) 

o Then tried to build a PLL. It locks easily. 
  - Put the beat signal into the mixer RF input.
  - Connect 10dBm @1MHz-10MHz from the marconi oscillator to the LO input. The supposed nominal LO level was 
not checked so far. Just used 10dBm.
  - The IF output was connected to an SR560 with 10Hz LPF (6dB/oct) with G=500 or so.We don't need to care 
about the sign.
  - The output of the SR560 was connected to FAST PZT input of the NPRO.
o The problem was that there was strong intermodulations because of 33MHz. No LPFwas used before the mixer. 
Because of this spourious modulations, the PLL servo locks at the local zero crossings. These will be solved 
next time. 

o Eventually left the 40m lab at 2AM. "Last autoalignment" was restored. The flipper for the inj beam was 
down. The shutter for the NPRO was closed.
  737   Thu Jul 24 21:53:00 2008 ranaSummaryTreasureHigh School Tour group and the PMC
There was a tour today of 40 high school kids. I warned them that the lasers could burn out their
eyes
, that the vacuum could suck them through the viewports like tubes of spaghetti and that the
high voltage amps would fry their hair off.

One of them was taking a picture of the SOS in the flow bench and another one was whispering what
a dumb idea it was to leave a sensitive clean optic out where people might breathe on it. I told
one them to cover his mouth. The other one asked what was the glass block behind the SOS.

It was a spare PMC! s/n 00-2677 with a 279 nF capacitance PZT. I guess that this is the one that
Go brought from MIT and then left here. So we don't have to take the one away from Bridge in the
35 W laser lab.

We can swap this one in in the morning while the FSS people work on the reference cavity
alignment. Please email me if you object to this operation.
  736   Thu Jul 24 21:04:58 2008 ranaUpdatePSLFSS
Since Jenne and Yoichi are going to finish up their refcav/FSS work in the morning I decided to
look at the trends. I set the RF modulation level from 10.0 back down to 7.5 so that we would
have the same RF modulation depth as before. I also set the FSS common gain and its nominal to
1.0 dB since it seemed more stable this way.

With 7.5, the transmission of the refcav is ~6.9 V. It was around 0.7 V before so there's already
been a factor of 10 improvement in the power since the work started. In addition to the mode matching
work which is about to commence, we should attenuate the RC TRANS with a real mirror (not ND) so that
the camera and PD don't saturate. We should also do the same for the REFL PD and camera and make sure
to put in a steering mirror for the REFL PD and orient REFL so that it faces West (so that we can
look at its face with a viewer) and dumps its reflection.

Since the common gain is so low now, I expect that we will want less light in total. We can achieve
this by turning down the RF drive to the VCO.

I also fixed the MC down script which was putting the FSS common gain to the unstable +10 dB level
during the MC locking process.
  735   Thu Jul 24 19:29:26 2008 YoichiConfigurationPSLC1:PSL-STAT_FSS_NOM_C_GAIN is changed from 30 to -0.7
Koji, Yoichi

Since the light power going to the ref. cavity is now significantly increased (see Janne's elog later),
C1:PSL-STAT_FSS_NOM_C_GAIN
is changed from 30 to -0.7.
Otherwise, the MC did not lock.
  734   Thu Jul 24 11:49:07 2008 MashaSummaryAuxiliary lockingbelated weekly summary
I designed a high pass filter to whiten the spectrum from the Mach Zehnder to optimize the
input into the ADC. The swept sine response measurement and the effect of the filter on the
spectrum are attached. If I start using the digital system (it is currently down in Bridge),
I will decide if the filter needs to be improved/better matched to the ADC there.

I moved from the 40m to Rana's lab in bridge. I am making a new and improved Mach Zehnder
setup with a 50m fiber in one arm; currently the transmission through the fiber is 44%. I
am working out how to mode match the laser to the fiber to improve this number.
Attachment 1: filter_tr_function.pdf
filter_tr_function.pdf
Attachment 2: filtered_spectr0724.pdf
filtered_spectr0724.pdf
  733   Thu Jul 24 08:09:26 2008 YoichiUpdateLSCArm cavity g-factor measurement

Quote:

A-ha! Do you always expect the 30Hz signal, don't you?
Because this is the PDH technique.


Yes you are right. I realized this when I was thinking about it in the bed Smile
The 30Hz signal should always be present because the carrier is phase shifted at 30Hz by the cavity length change.
I think the change in the signal ratio between I and Q happened because as the 166MHz sidebands get phase change when they move around the MC transmission peak due to the cavity pole of the MC. It changes the optimal demodulation phase for the 166MHz PDH signal at the AS port.


Quote:

We should be able to see 166MHz sideband resonances using the double demodulated photodetectors. With these, the 33MHz sidebands will be acting as LO when the 166MHz sideband (or mode) resonates. Some modeling may be necessary to determine if the SNR will be good enough to make this worthwhile, however.


I will try, but at 100kHz away from the MC FSR (the number predicted by John's calculation), the transmission of the 166MHz sidebands is very weak. I did not see any signal when I swept it +/- 500kHz. Unfortunately, the Marconi's output level is almost at its maximum. So we don't have much room for increasing the sideband power.
  732   Thu Jul 24 03:08:20 2008 robUpdateLocking+f2 DRMI+2ARMS

rob, john, yoichi

Tonight we tried to move the 166MHz (f2) sideband frequency by changing the settings on the Marconi. Reducing the frequency by 4kHz reduced the amplitude of the 166MHz sidebands, but we were still able to lock the DRMI with the +-f2 sidebands by electronically compensating for the gain decrease, and also to lock the DRMI+2ARMs while resonating the -f2 sideband. No luck with the +f2.

Then we larkily tried increasing the frequency by 4kHz, which ~doubled the f2 sideband transmission through the MC. This means our frequencies/MC length have been mismatched for months. Apparently I explained the level of the f2 sidebands by just imagining that I'd (or someone) had set the modulation depth at that level some time in the past.

It's a miracle any locking worked at all in this state. Once this was done and we worked out a few kinks in the script, adjusting some gains to compensate, we managed to get the DRMI+2ARMS to lock a couple of times while resonating the +f2 sideband. It takes a while, but at least it happens. Tomorrow we'll measure the length of the mode cleaner properly and then try again. No need to vent just yet.
  731   Thu Jul 24 02:57:26 2008 robUpdateLSCArm cavity g-factor measurement

Quote:

So, now I feel that the method for the TEM01 quest should be reconsidered.

If we have any unbalanced resonance for the phase modulation sidebands, the offset of the error signal is to be observed even with the carrier exactly at the resonance. We don't need to shake or move the cavity mirrors.

Presence of the MC makes the things more complicated. Changing the frequency of the modulation that should go throgh the MC is a bit tricky as the detuning produces FM-AM conversion. i.e. The beam incident on the arm cavity may be not only phase modulated but also amplitude modulated. This makes the measurement of the offset described above difficult.

The setup of the abs length measurement (FSR measurement) will be easily used for the measurement of the transverse mode spacings. But it needs some more time to be realized.


We should be able to see 166MHz sideband resonances using the double demodulated photodetectors. With these, the 33MHz sidebands will be acting as LO when the 166MHz sideband (or mode) resonates. Some modeling may be necessary to determine if the SNR will be good enough to make this worthwhile, however.
  730   Thu Jul 24 01:27:00 2008 KojiUpdateLSCArm cavity g-factor measurement

Quote:
I locked the Y-arm and mis-aligned the end mirror in Yaw until the transmission power gets half.
Then I injected a 30Hz sinusoid into the error point of the Y-arm servo to shake the ETMY.
I observed AS166_I and AS166_Q as I changed the 166MHz frequency.


A-ha! Do you always expect the 30Hz signal, don't you?
Because this is the PDH technique.

---------------
Recipe:
You have a carrier and phase modulation sidebands at 166MHz this time.
Inject them into a cavity. Detect the reflection by a photo detector.
Demodulate the photocurrent at 166MHz.

This is the PDH technique.

A 30Hz sinusoid was injected to the error point of the cavity lock.
This means that the cavity length was fluctuated at 30Hz.

We should see the 30Hz signal at the error signal of the 166MHz demodulation, regardless of the tuning of the modulation frequency!
In other words, the 30Hz signal in the demod signal at the 166MHz is also understandable as the beating between the 30Hz sidebands and the 166MHz sidebands.

---------------

So, now I feel that the method for the TEM01 quest should be reconsidered.

If we have any unbalanced resonance for the phase modulation sidebands, the offset of the error signal is to be observed even with the carrier exactly at the resonance. We don't need to shake or move the cavity mirrors.

Presence of the MC makes the things more complicated. Changing the frequency of the modulation that should go throgh the MC is a bit tricky as the detuning produces FM-AM conversion. i.e. The beam incident on the arm cavity may be not only phase modulated but also amplitude modulated. This makes the measurement of the offset described above difficult.

The setup of the abs length measurement (FSR measurement) will be easily used for the measurement of the transverse mode spacings. But it needs some more time to be realized.
  729   Thu Jul 24 01:04:01 2008 robConfigurationLSCIFR2023A (aka MARCONI) settings

Quote:


P.S.: We made a test by changing the frequency of the local oscillator by a little bit and then coming back to the original value. We observed that the phase of the signal can change, so every time this frequency is moved the 3f demod phase need to be retuned.



We discovered this little tidbit in March, and remembered it tonight. Basically we found that whenever you change the frequency on one of these signal generators (and maybe any other setting as well), the phase of the signal can change (it's probably just the sign, but still...), meaning that you when you return settings to their intial value, not everything is exactly as it once was. For most applications, this doesn't matter. For us, where we use one Marconi to demodulate the product of two other Marconis, it means we can easily cause a great deal of grief for ourselves, as the demod phase for the double demod signals can appear to change.

Programmatically, what this means is that every time you touch a Marconi you must elog it. Especially if you change a setting and then put it back.
  728   Wed Jul 23 22:34:07 2008 YoichiUpdateLSCArm cavity g-factor measurement
I tried the same thing as the X-arm to the Y-arm.
I'm puzzled. I found exactly the same behavior as the X-arm in the AS166 demodulated signals, whereas I expected different resonance frequency because of the arm length difference.

Here is more detailed account of the measurement today.

I locked the Y-arm and mis-aligned the end mirror in Yaw until the transmission power gets half.
Then I injected a 30Hz sinusoid into the error point of the Y-arm servo to shake the ETMY.
I observed AS166_I and AS166_Q as I changed the 166MHz frequency.

At 165.977MHz, both AS166_I and AS166_Q showed the 30Hz signal (15cnt p-p).
At 165.981MHz, Only I phase showed the 30Hz signal (40cnt p-p). No signal in Q.
At 165.984MHz, I and Q became the same amplitude again (20cnt p-p).
At 165.987MHz, Only Q phase showed the 30Hz signal (40cnt p-p). No signal in I.

Outside the above range, the signal decreases as the frequency go away. I think this is (at least partly) because the 166MHz sidebands no longer go through the MC at those frequencies.

I then locked the X-arm to the TEM01 mode. I saw exactly the same behavior as described above. This could be the resonance of TEM02 mode. I was expecting to see the resonance of TEM00 mode at the opposite side, but nothing there.

I unlocked the arm cavities and tried the same frequency scan of the 166MHz with one of the end mirrors shaken at 30Hz. I saw no signal at the AS166 port.
I also tried locking Y-arm and shaking the ETMX. No signal.
So it has to be something to do with the cavity resonance.

Since the MC transmission curve for 166MHz is folded in the measurement, it makes the interpretation of the results harder.
  727   Wed Jul 23 21:48:30 2008 robConfigurationGeneralrestore IFO when you're done with it

when you are done with the IFO, please click "Restore last auto-alignment" on the yellow IFO portion of the C1IFO_CONFIGURE.adl screen. Failure to comply will be interpreted as antagonism toward the lock acquisition effort and will be met with excoriation.
ELOG V3.1.3-